OpenStudy (hihi):

If the linear equation 2x = 9 − 3y is written in standard form, what is the value of B?

OpenStudy (karatechopper):

well standard form is Ax+By=C so u want to first get it in that form

OpenStudy (karatechopper):

so can you write your equation u have for me in standard form?

OpenStudy (hihi):

sure

OpenStudy (hihi):

2x+3y=9

OpenStudy (hihi):

wait i didnt change the signs...

OpenStudy (karatechopper):

yes u did its correct

OpenStudy (hihi):

oh my bad thank you

OpenStudy (hihi):

so the value of b would be 3? Because i think there is more.

OpenStudy (karatechopper):

haha its fine:) so u want to find B. well B is the y intercept. so can u tell me what the y intercept in this equation?

OpenStudy (hihi):

oh ok

OpenStudy (karatechopper):

i meant 3 soorry

OpenStudy (karatechopper):

the y intercept is not 3

OpenStudy (hihi):

would the y intercept be 3 because its paired with b?

OpenStudy (karatechopper):

no

OpenStudy (karatechopper):

lets start from here: y=mx + b can u tell me which form this equation is written in?^

OpenStudy (hihi):

slope intercept form

OpenStudy (karatechopper):

correct! and can u tell me what each variable stands for?

OpenStudy (hihi):

so the slope stands for mx right?

OpenStudy (hihi):

but the slope is what we are trying to find.

OpenStudy (hihi):

im so confused right now

OpenStudy (karatechopper):

woah there! look... this is what the variables stand for.. m=slope b=y intercept y= what u solved for x= intercept

OpenStudy (hihi):

ok ya i remember

OpenStudy (karatechopper):

notice that in the equation of slope intercept form.. y=mx+b b has no variable connected with it.

OpenStudy (karatechopper):

so in your equation.. 2x+3y=9 which number has no variable connected to it?

OpenStudy (hihi):

ok

OpenStudy (hihi):

9

OpenStudy (karatechopper):

correct!!

OpenStudy (hihi):

so what would the answer be?

OpenStudy (karatechopper):

9 because it is in place of the letter b

OpenStudy (hihi):

thank you :)

OpenStudy (karatechopper):

no problemo:)